A particle is projected with a velocity of 64m/s at angle 45 degree to the horizontal. find the highest height attained by the particle..g= 10

Grade:9

1 Answers

Vikas TU
7 years ago
Maximum Height of a projectile is given as:
H = u^2sin^2(thetha)/g
but to be maximum thetha should be 90 degree.
And therefore,
H = 64*64/2g = > 204.8 meter.